OpenStudy (anonymous):

Solve for x: 3x-1/5-2x-4/3=1 -38 38 -2 2

OpenStudy (anonymous):

\[\frac{ 3x-1 }{ 5 }-\frac{ 2x-4 }{ 3 }=1 \]

OpenStudy (anonymous):

@Yttrium

OpenStudy (yttrium):

Just multiply the whole equation by 15. :)